OpenStudy (sleepyjess):

Mandy told Bill that she has read 16 books this year and reads 2 books each month. Bill wants to catch up to Mandy. He tracks his book reading with a table on his door. Using his table below, what month will Bill have read the same amount of books as Mandy? Month Books May 4 June 8 July 12 I know the answer is November, but I don't understand how to get to that answer. What formula would I use?

OpenStudy (anonymous):

Let's see, So mandy has already read 16 books and she reads 2 each month That would be: 16+2x ( x=months) For Bill now. He hasnt read any yet but reads 4 books per month so: 4x Let's equate these and we get 16+2x=4x 2x = 16 x = 8 so in 8 months bill will have read the same books as mandy which is 32 books :)

OpenStudy (sleepyjess):

The answer was November though.

jigglypuff314 (jigglypuff314):

sorry for taking so long :/ It was phrased slightly odd... because it hadn't given Mandy's starting point... this was what I tried: x = 0 1 2 3 4 5 6 May June July Aug Sept Oct Nov M: 16 18 20 22 24 26 28 B: 4 8 12 16 20 24 28 so formula-wise you would get M = 2x + 16 B = 4x + 4

OpenStudy (anonymous):

A little off topic here but does this remind anyone of the cartoon Billy and Mandy?

OpenStudy (anonymous):

The question didn't state anything about when mandy or bill started so i assumed that they start from the beggining of the year :S

jigglypuff314 (jigglypuff314):

I was getting December at first as well, I managed to get November in the end... the confusing part was whether Mandy had started with reading 16 books in April or May if she had 16 in April then it would be x = 8 aka December ("16+2x=4x") if she had 16 in May then it would be x = 6 aka November (16 + 2x = 4x + 4)

OpenStudy (sleepyjess):

And that was where my confusion came in. They really need to be more specific/give more detail on these problems! lol

OpenStudy (anonymous):

@jigglypuff314 yeah thats the only way you could get november so that must be it! it's still odd tho because if you assume that we are on august from bill's table the other clue that we get about mandys 16 books is from may! it should say that mandy has read 20 books not 16!

jigglypuff314 (jigglypuff314):

I hope this helped @sleepyjess sorry for taking so long, and thank you for your patience on this problem :)

OpenStudy (sleepyjess):

Thank you!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

i worked it out and it wasn't the December, the correct answer is November. i just took the test

jigglypuff314 (jigglypuff314):

@henryarias5 yes indeed :) we managed to work it out, if you read one of my previous comment: "I was getting December at first as well, I managed to get November in the end... the confusing part was whether Mandy had started with reading 16 books in April or May if she had 16 in April then it would be x = 8 aka December ("16+2x=4x") if she had 16 in May then it would be x = 6 aka November (16 + 2x = 4x + 4)"

OpenStudy (anonymous):

i got the answer right it's December for sure for sure and i think @ocelot40 is right when she said Let's equate these and we get 16+2x=4x 2x = 16 x = 8 so in 8 months bill will have read the same books as mandy which is 32 books count 8 from may u will get December , they have updated the answer i'm assuming because i put November at first and i got it wrong and when i retook the exam i got December as the right answer

jigglypuff314 (jigglypuff314):

^ they probably realized that they worded the question wrong and so changed it properly xD so a note to everyone looking to cheat and came to this question post - congrats you've read this notice. NOW STOP CHEATING AND GO AWAY :3 [yes I do realize when people somehow medal me on a year old post >.>]

OpenStudy (sleepyjess):

LOL

OpenStudy (coollikedat):

Just strolling by.... ;)

jigglypuff314 (jigglypuff314):

yah know... if you want to get away with cheating... a good thing to do is *not* alert the person that has the ability to ban people that you are cheating ^_^ [cough] I can suspend people I do get notifications every time that someone medals me Please be more careful xD